International Congress of Mathematicians Beijing, China August 20-28, 2002

Second Announcement

The Organizing Committee is pleased to invite you to attend the International Congress of Mathematicians in Beijing, August 20–28, 2002.

Mr. Jiang Zemin, the President of the People’s Republic of China, has personally expressed a most cordial welcome to all participants.

The Congress will be held under the auspices of the International Mathematical Union and under the sponsorship of the Chinese Mathematical Society (CMS), the Ministry of Education, the Ministry of Science and Technology, Chinese Academy of Sciences and the National Natural Science Foundation of China.

This announcement describes the Congress and gives related information. It explains how to register and how to submit a short communication or a poster representation. It also contains the necessary forms for securing accommodation during the Congress.

Contents

A. Location of the Congress 1 A.1 Venue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 A.2 The Congress . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 A.3 Deadlines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

B. Scientific Program 2 B.1 Opening and Closing Ceremonies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 B.2 Plenary Lectures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 B.3 Invited Lectures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 B.4 Short Communications and Posters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 B.5 Organized Sessions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 B.6 Informal Seminars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

C. Other Events of Mathematical Interest 5 C.1 Events for the General Public . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C.2 Book Exhibitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C.3 Other Mathematical Conferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

D. Publications 6 D.1 Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 D.2 List of Participants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 D.3 Abstracts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 D.4 The Beijing Intelligencer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 D.5 Proceedings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 D.6 Daily News . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 E. Social Program 8

E.1 Opening Reception . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 E.2 ICM2002 Party . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 E.3 Beijing Opera . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 E.4 Tourist Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 E.5 Footloose Tours . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 E.6 Activities for Accompanying Persons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

F. Travel 9 F.1 Passports and Visas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 F.2 Congress Agent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 F.3 Arrival in Beijing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 F.4 Local Transportation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

G. Mail and Messages 10 G.1 Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 G.2 Telephone Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 G.3 Personal Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 G.4 Public E-Mail Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

H. Miscellaneous Information 11 H.1 Official Languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 H.2 Invitation Letter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 H.3 Climate and Clothing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 H.4 Electric Current . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 H.5 Bank Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 H.6 Credit Cards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 H.7 Shopping Hours . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 H.8 First Aid, Health and Accident Insurances . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

I. Registration 12 I.1 Membership . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 I.2 Accompanying Persons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 I.3 Registration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 I.4 Registration Fees . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 I.5 Methods of Payment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 I.6 Cancellations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 I.7 Registration Counter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

J. Accommodation 14 J.1 Hotels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 J.2 Apartments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 J.3 Student Dormitories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 J.4 Liability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

K. Social and Tourist Program 16 K.1 Social Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 K.2 Tourist Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

L. Satellite Conferences 20

M.ICM2002 Sponsors 23

N. 2000 Mathematics Subject Classification 24

O. Organizing Committee 25

A. Location of the Congress

A.1 Venue

Mathematical traditions in Beijing can be traced back to ancient times. In the development of mathematics in modern China, Beijing has played an increasingly important role. In the early 1930’s, a group of outstanding Chinese mathematicians, including Hua Luo-keng and Chern Shiing-Shen, were trained here and stepped on to the international stage of mathematics from here. Today, Beijing is one of China’s main centers for mathematical research and education. Three leading universities (Peking University, Tsinghua University, and Beijing Normal University) and four research institutes of the Academy of Mathematics and System Sciences of the Chinese Academy of Sciences (Institute of Mathematics, Institute of Applied Mathematics, Institute of Systems Science, and Institute of Computational Mathematics and Scientific/Engineering Computing), as well as the Nankai Institute of Mathematics at Nankai University in the nearby city of Tianjin, conduct mathematics instruction on a broad front and carry out many state-of-the-art research projects in various fields. In recent years, special attention has been paid to attracting and training excellent young mathematicians, and international exchanges have greatly increased. Every year we receive numerous visitors from all over the world.

Beijing is the capital of the People’s Republic of China. It is China’s political, cultural, educational and research center. With a history of 3000 years, it has been the nation’s capital for the last 800 years and thus has much to offer to visitors, such as the Great Wall, the Forbidden City, the Temple of Heaven, the Summer Palace, etc.

A.2 The Congress

The activities of the Congress are divided between two locations. The program of the first day, August 20, 2002, including the opening ceremony, will take place at the Great Hall of the People. As the Great Hall is located in the center of the city, transportation will be provided to delegates staying at the congress-assigned hotels. On all other days, the lectures, seminars and short communications will be held at the Beijing International Convention Center (BICC).

The address of the Congress is:

ICM2002

c/o Prof. Ya-xiang Yuan

Academy of Mathematics and System Sciences Chinese Academy of Sciences 1A South 4th Street Zhong Guan Cun, Beijing 100080 People’s Republic of China

Phone: +86 10 6256-0649 Fax: +86 10 6261-8223

     E-mail: icmsec@beijing.icm2002.org.cn
     WWW: http://www.icm2002.org.cn

A.3 Deadlines

April 30, 2002 Early Registration at reduced rate
May 1, 2002 Submission of abstracts for

• Plenary Lectures

• Invited Lectures

• Short Communications

• Poster Sessions

May 15, 2002 Apartment reservation
June 15, 2002 Hotel reservation
June 15, 2002 Submission of manuscripts of

• Invited Lectures for the Proceedings • Plenary Lectures for the Proceedings

July 15, 2002 Cancellation (with partial refund) of

• registration

B. Scientific Program

The final program will be available electronically by June 1, 2002.

B.1 Opening and Closing Ceremonies

The opening ceremony will be held in the Great Hall of the People at 15:00 on Tuesday, August 20, 2002. As part of the program, the Fields medals and the Nevanlinna prize will be awarded. The Great Hall of the People seats up to 10000 people and allows a good view from all seats.

After the opening session, there will be addresses on the works of the Fields medalists and the Nevanlinna prize winner.

The closing ceremony, after the last two Plenary Lectures, is scheduled for Wednesday, August 28, at 16:00 in the main lecture hall of BICC.

B.2 Plenary Lectures

At the recommendation of the Program Committee, appointed by the International Mathematical Union (IMU), the Organizing Committee has invited 20 mathematicians to give one-hour Plenary Lectures. The names of the speakers, together with their affiliations and fields of research, are listed below. The Plenary Lectures will inform participants of major developments, problems, and trends in mathematics.

The IMU General Assembly, the Program Committee, and the Organizing Committee have emphasized that these lectures should be comprehensible to a wide spectrum of mathematicians. All plenary speakers have agreed to prepare addresses for a general mathematical audience.

Noga Alon (Tel Aviv Univ., Israel): Discrete Mathematics Douglas N. Arnold (Univ. of Minnesota, USA): Numerical Analysis, Differential Equations,

Mechanics, Computational Relativity

Alberto Bressan (S.I.S.S.A. Italy): Hyperbolic Systems of Conservation Laws, Nonlinear Control

Theory

Luis Angel Caffarelli (Univ. of Texas at Austin, USA): Partial Differential Equations Sun-Yung Alice Chang (Princeton Univ., USA): Geometric Analysis David Leigh Donoho (Stanford Univ., USA): Statistics Ludwig D. Faddeev (St. Petersburg Dept, Steklov Math. Inst., Russia): Mathematical Physics Shafi Goldwasser (MIT, USA and Weizmann, Israel): Computer Science Uffe Haagerup (Univ. of Southern Denmark, Denmark): Operator Algebras Michael Jerome Hopkins (MIT, USA): Algebraic Topology Victor G. Kac (MIT, USA): Algebra and Mathematical Physics Harry Kesten (Cornell Univ., USA): Probability Theory Frances Clare Kirwan (Univ. of Oxford, UK): Symplectic Geometry, Algebraic Geometry Laurent Lafforgue (IHES, France): Arithmetic Algebraic Geometry and Automorphic

Representations

David Mumford (Brown Univ., USA): Statistics, Applied Mathematics Hiraku Nakajima (Kyoto Univ., Japan): Geometry, Representation Theory Yum-Tong Siu (Harvard Univ., USA): Several Complex Variables, Differential Geometry,

Algebraic Geometry

Richard Lawrence Taylor (Harvard Univ., USA): Number Theory Gang Tian (MIT, USA and Peking Univ., China): Geometry, Nonlinear Analysis Edward Witten (Princeton Inst. for Adv. Study, USA): Quantum Field Theory, String Theory

All the addresses will be given in Hall 1, the main lecture hall, of BICC, which has a capacity of 2500 seats. If necessary, the talks will also be shown in an adjoining lecture hall via closed-circuit television, so that another audience of more than 1000 people can be reached.

B.3 Invited Lectures

Also at the recommendation of the IMU Program Committee, 169 mathematicians have been invited to give 45-minute lectures in specified sections. These lectures are intended to be surveys of significant topics in the respective area of research. The speakers have been asked to make their lectures comprehensible for the general mathematical community, and they agreed to do so.

The Invited Lectures will be given in several large lecture rooms at BICC. Usually there will be about six or seven lectures simultaneously.

The list of sections is as follows, where the number in parentheses indicates the expected number of presentations in that section.

1. Logic (5)
2. Algebra (9)
3. Number Theory (10)
4. Differential Geometry (14)
5. Topology (9)
6. Algebraic and Complex Geometry (8)
7. Lie Groups and Representation Theory (11)
8. Real and Complex Analysis (9)
9. Operator Algebras and

Functional Analysis (6)

10. Probability and Statistics (13)
11. Partial Differential Equations (12)
12. Ordinary Differential Equations and

Dynamical Systems (11)

13. Mathematical Physics (12)
14. Combinatorics (8)
15. Mathematical Aspects of Computer Science (7)
16. Numerical Analysis and

Scientific Computing (7)

17. Applications of Mathematics in the Sciences (12)
18. Mathematics Education and

Popularization of Mathematics (3)

19. History of Mathematics (3)

In addition, 8 mathematicians have been invited to give talks in two panel discussion sessions of Section 18. Each panel discussion session lasts 90 minutes with 4 invited speakers.

B.4 Short Communications and Posters

All Ordinary Members (see I.1) of the Congress will have the opportunity to present their mathematical work in the form of a Short Communication or a Poster — provided that

• they have registered by May 1 and marked in the registration form, that they want to present their work,

• they have submitted an abstract by that date,

• their contribution has been accepted by the Local Scientific Committee.

Only one Short Communication or Poster (and thus only one abstract) is allowed for each member. Each Short Communication lasts 15 minutes including discussion. Short Communications are grouped into time slots of 45 minutes for three presentations. The rooms for Short Communications are equipped with an overhead projector. Each Poster Session lasts 105 minutes; during that period the authors should stand by their posters and be available for questions and discussion.

Authors presenting a poster are advised to bring the material of the Poster with them when they come to the Congress since no facilities for preparing posters are available on site. The size of the individual poster panels is as follows: width 100 cm, height 245 cm.

The abstract for a Short Communication or a Poster must include the appropriate section number (see B.3) and 2000 MS classification number (see N.) so that the Communications and Posters can be grouped in a coherent way for presentation. Abstracts may be submitted in English, French, Russian or German. Abstracts of accepted Short Communications and Posters which are properly prepared and received by the deadline will be reproduced and distributed to all Ordinary Members when they pick up their registration package. The Local Scientific Committee will notify authors of acceptance/rejection of their contribution.

Instructions on how to prepare an abstract are in Section D.3. Abstracts which do not conform to the stipulated rules will be returned to the author for resubmission. Late papers will not be accepted. However, it is possible to present them in ad-hoc sessions that will be organized and announced during the Congress.

B.5 Organized Sessions

We invite Ordinary Members to organize sessions of Short Communications on their own initiative. Such sessions and their organizers will be included in the final program after a review and consent by the Local Scientific Committee. Ordinary Members who want to organize such a session should contact the Local Scientific Committee at scipcomm@beijing.icm2002.org.cn with their program and their request for time slots within the schedule of Short Communications.

All participants of these organized sessions have to submit an abstract for their Short Communication as indicated in Section D.3.

B.6 Informal Seminars

During the Congress it will also be possible to organize informal mathematical seminars on site.

Ordinary Members who wish to organize such spontaneous seminars are asked to make all arrangements among themselves, and to request a room either in advance (scipcomm@beijing.icm2002.org.cn), or during ICM2002 from the congress office. Such seminars may take place during the afternoons. If the congress office is notified before 15:00 the previous day, an announcement of the seminar can be included in the daily newsletter to be distributed to all participants on the following day.

C. Other Events of Mathematical Interest

C.1 Events for the General Public

A small number of lectures of non-technical nature will be arranged. These lectures are intended for the general public and they will highlight the important ways in which mathematics and mathematical ways of thinking are having an important impact on everyday life. The International Congress provides also other opportunities to inform the general public about some of the recent developments and future challenges of mathematics.

C.2 Book Exhibitions

Book, educational media, and computer software exhibits are located in the same building as the Congress, and are open to all Ordinary Members and accompanying persons during the Congress, from 8:00 to 18:00 on August 21 to 24 and 26-28, 2002,

C.3 Other Mathematical Conferences

Traditionally there are several smaller conferences scheduled at various places immediately before or after the Congress. On page 20, you can find a list of such satellite conferences (as of December 1, 2001), together with the addresses where further information may be obtained. You may also find further information at the website http://www.icm2002.org.cn. Please do not direct inquiries about these conferences to ICM2002.

D. Publications

The material described in sections D.1–D.4 will be included in the registration package that each Ordinary Member will receive at the registration counter.

D.1 Program

All Ordinary Members (see I.1) will receive copies of the official ICM2002 program. The program will show in detail the dates, times and locations of all Plenary and Invited Lectures. It will also contain a complete listing of all Short Communications and Poster presentations. See the ICM2002 WWW server http://www.icm2002.org.cn, after June 1, 2002.

D.2 List of Participants

A list of all participants who have registered by August 1, 2002 will be distributed to all Ordinary Members. A list of all ICM2002 members, including their mailing addresses, will be kept at the ICM2002 registration counter throughout the Congress. Ordinary Members are asked to check their own listing for accuracy while they are in Beijing, as this list will be used to prepare the official list of participants for inclusion in the Proceedings and to prepare mailing labels for shipment of the Proceedings.

D.3 Abstracts

Abstracts of Plenary and Invited Lectures, Short Communications and Posters will be reproduced and distributed in printed form to all Ordinary Members at the beginning of the Congress. These abstracts will also be available on the internet under the following addresses:

   http://www.mathca.com/icm2002.html (Plenary and Invited Lectures)

http://www.icm2002.org.cn/D/Abstract.htm (Short Communications and Posters)

Abstracts of Short Communications and Posters should be written in English, French, Russian or German and should have the following form (compare also the enclosed example):

– Section Number (see B.3)

– 2000 Mathematics Subject Classification number (see N.)

– Name and affiliation of author(s)

– Title

– Abstract text (no more than 120 words)

Example Abstract

Section: 2

2000 MS Classification: 17, 18, 55

Loday, Jean-Louis, Université de Strasbourg, France: Leibniz algebras and their (co)homology. A Leibniz algebra is a vector space equipped with a product satisfying a variation of the Jacobi identity: [x, [y, z]] = [[x, y], z] − [[x, z], y]. There is a dual notion in the sense of Koszul duality for operads. For any Leibniz algebra g there is a (co)homology theory HL(g), which satisfies various properties including the following: HL^{∗}(g) is a dual Leibniz algebra. Applications to non-commutative rational homotopy theory will be presented. Part of these results is joint work with T. Pirashvili. Reference: J.-L. Loday, and T. Pirashvili, Universal enveloping algebras of Leibniz algebras and (co)homology, Math. Ann. 296 (1993), 139–158.

Abstracts of Plenary Lectures and Invited Lectures should be written in the same way except that they should be written in English and the length of the body text should not exceed 600 words. All abstracts should be submitted electronically via the forms provided by one of the following WWW addresses:

http://www.mathca.com/icm2002.html (Plenary and Invited Lectures)
http://www.icm2002.org.cn/D/Abstract.htm (Short Communications and Posters)

or by e-mail using the subject “ICM2002 abstract” to one of the following e-mail addresses:

icm2002@fudan.edu.cn (Plenary and Invited Lectures) editor@beijing.icm2002.org.cn (Short Communications and Posters)

TEX and html forms supporting electronic submission will be available under the above WWW address.

Submission is also possible by fax or by ordinary mail (preferably with the file in a 3.5-inch diskette) to

Plenary and Invited Lectures:

ICM2002 Abstracts

c/o Prof. Li Ta-tsien

Department of Mathematics Fudan University 220 Han Dan Road Shanghai 200433, China Fax: +86 21 6564-8274

Short Communications and Posters:

ICM2002 Abstracts c/o Prof. Ding Weiyue Academy of Mathematics and System Sciences 1A South 4th Street Zhong Guan Cun, Beijing 100080, China Fax: +86 10 6261-8223

However, electronic submission at the WWW web server is strongly encouraged. The deadline for submission of abstracts is May 1, 2002.

D.4 The Beijing Intelligencer

The Chinese Mathematical Society, Springer-Verlag Berlin/Heidelberg and Higher Education Press will together prepare, publish, and present to all participants a magazine called the “Beijing Intelligencer”, which is meant to welcome you to Beijing and to the ICM2002 and to guide your visit to Beijing and to some of its manifold aspects and attractions.

D.5 Proceedings

All Plenary and Invited Lectures as well as the congress report will be published in the Proceedings of the International Congress of Mathematicians, Beijing, 2002. These Proceedings will be published by Higher Education Press.

A printed version of the proceedings will be distributed to all Ordinary Members and students of the Congress. There will also be an electronic version published in the International Math Arxiv at the website: http://front.math.ucdavis.edu/ICM2002/ or http://www.cgtp.duke.edu/ICM2002/.

D.6 Daily News

A newsletter containing program changes, announcements of informal seminars and information of general interest to ICM2002 participants will be available each day outside the meeting rooms in the BICC. Participants with announcements for the newsletter should be sure to submit them to the conference office no later than 15:00 the day before.

E. Social Program

E.1 Opening Reception

On August 20, an Opening Reception will be offered to all participants immediately after the Opening Session. The opening reception will be held at the Banquet Hall of the Great Hall of the People. It is free for all participants and registered accompanying persons.

E.2 ICM2002 Party

On the evening of August 26, an ICM2002 Party will take place for all participants and registered accompanying persons. If weather permits, it will be held in the open, on the lawn.

E.3 Beijing Opera

Beijing Opera is a purely Chinese operatic form with a history of over 200 years. It is a combination of singing, dancing, Kongfu, acrobatics, music, arts and literature.

On the evening of August 25, 2002, at 19:30, a show of Beijing Opera will be arranged at Chang’an Theater. The show contains three dramatic selections from the well-appreciated complete plays. For the time being, the three selections are The Crossroads, Stealing Magic Herbs, and Farewell My Concubine (The Death of Yu Ji). Those who are interested may make reservations with the attached form. Transportation will be provided to those staying in the congress-assigned hotels.

E.4 Tourist Program

As the heart of the nation from the height of China’s dynastic splendor 600 years ago and home to 24 emperors who ruled China’s vast territory, Beijing has been bequeathed with many historical sites of great aesthetic and cultural value. During the Congress, guided local tours will be organized to provide participants and accompanying persons an opportunity to appreciate the splendid culture of the Chinese nation. See Section K.2 for details.

E.5 Footloose Tours

Footloose tours will be offered as informal tours to special museums, parks and shopping streets. The tours are free of charge, except for possible entrance fees. These tours will be announced shortly before the Congress by means of Circular Letters from the ICM2002 e-mail service and posted on the WWW-server of ICM2002 http://www.icm2002.org.cn.

E.6 Activities for Accompanying Persons

Accompanying persons, whether registered or not, are welcome to participate in all activities of the social program. They are invited to take part in the informal Footloose Tours and the formal Local Tours. However, only registered accompanying persons can attend the Opening ceremony and the Opening Reception in the Great Hall of the People, and the ICM2002 Party.

F. Travel

F.1 Passports and Visas

Passports, valid for at least six months from the entry into China, and entry visas are required to visit China. Visas may not be required for visitors from certain countries, and Congress participants are advised to consult the nearest Chinese diplomatic mission for details. A single entry visa is valid for three months, and can be extended for an additional month at the Foreigners Section of the Beijing Public Security Bureau if necessary. Once in China, the Congress participants are advised to carry their passports on them at all times because the passports may be needed to check into hotels, book airline tickets, and change money. To apply for visas, Congress participants should first fill out and send in the registration form (see attached) with the appropriate registration fee. The Congress Secretariat will then send a formal invitation via airmail with which the Congress participants can apply for a visa at the Chinese diplomatic mission indicated in the invitation. This application should be made preferably at least one month before departure for China to allow time for processing. Alternatively, Congress participants may ask their travel agents to arrange their travel to China and obtain a tourist visa irrespective of whether or not they have registered with the Congress Secretariat.

F.2 Congress Agent

China International Conference Center for Science and Technology (CICCST) has been appointed by the Organizing Committee to handle registration for the Congress, and reservation of accommodation and tours, etc. for the Congress participants. Please send all correspondence related to the scientific program of the Congress to the official congress address as indicated in A2, and that related to registration and reservation of hotels and tours to the CICCST at:

ICM2002 Secretariat Attn. Mr. Liu Feng CICCST 86 Xueyuan Nan Road Beijing 100081, China Tel: +86 10 6217-4952 Fax: +86 10 6218-0142

     Email: icmagent@beijing.icm2002.org.cn

F.3 Arrival in Beijing

Beijing can be reached by air or by train. Capital International Airport is the only airport for civil flights. If you arrive by air, you should better first exchange some Chinese money at the Bank of China Airport Service in the Terminal Building, so as to be able to pay for taxi and other minor costs. Then you may easily get a taxi at the taxi stand outside the airport terminal. Make sure that the taximeter is turned on and you get a receipt printed from the taximeter. The cost for a taxi from the airport to the Congress venue is roughly 100 yuan or 13 US dollars, including the highway toll fee. There are regular shuttle buses to town, but they do not pass your hotel or the Congress venue. Please be advised that you are not supposed to accept help of taxi service from anyone other than at the taxi stand.

The Organizing Committee will send some staff, with the sign of the Congress logo, to the airport on August 18, 19 and the moroning of 20th to meet the arriving Congress participants and to direct them. In case you cannot get their help, the following message, written in Chinese, may help you get to the Beijing International Convention Center:

The meaning of the above Chinese message is “please take me to Beijing International Convention Center in the Asian Games Village”. A sketch map of Beijing and that of the neighborhood of BICC are given in pages 26 and 27 respectively.

If you arrive by train, take the subway train to Anding Men, and then take Bus No. 406 or taxi to Ya Yun Cun (Beijing International Convention Center).

F.4 Local Transportation

Beijing has a perfect public bus network by which you can get almost anywhere, but taxi is the most convenient for foreign visitors.

The Organizing Committee has arranged transportation between the Congress venue and the designated hotels of Category B, C, D(see J.1) and student dormitories(see J.3). Each morning, buses will take you to the venue; in the late afternoon, they will take you back to your hotel.

G. Mail and Messages

G.1 Mail

All mail, telegrams, and faxes for persons attending the Congress should be addressed to:

Name of the Participant

c/o ICM2002 , Prof. Ya-xiang Yuan

Academy of Mathematics and System Sciences Chinese Academy of Sciences 1A South 4th Street, Zhong Guan Cun Beijing 100080, China Tel: +86 10 6256-0649 Fax: +86 10 6261-8223

Incoming items will be posted at the conference office in the BICC.

G.2 Telephone Messages

For urgent messages to a participant, call

+86 10 6256-0649

or send a fax to

+86 10 6261-8223

To avoid mistakes, only simple messages will be taken and will be put in the respective mailbox. Unless it is a very urgent matter, it will not be possible for our secretarial staff either to check the presence of a specific person or to check whether the message actually reaches the intended person.

Participants can make international telephone calls from their hotel rooms, or at a public telephone booth with a phone card, which is sold at the General Service Counter in the Beijing International Convention Center and in the post office.

G.3 Personal Messages

Participants wishing to exchange personal messages during ICM2002 should stick their message on the Message Board erected in the lobby of BICC.

G.4 Public E-Mail Service

During the congress days public e-mail service will be available to Ordinary Members of ICM2002. For this, a mail server and some client computers are offered in the BICC. Opening hours during the Congress are from 9:00 to 20:00.

H. Miscellaneous Information

H.1 Official Languages

English, French, Russian, and German are the official languages of the Congress. Announcements, correspondence, and all other business matters will be carried out in English.

H.2 Invitation Letter

An Official Invitation Letter will be sent by the Organizing Committee upon request (see the second page of the registration form). This personal invitation is intended only to facilitate participants’ travel and visa arrangements.

H.3 Climate and Clothing

Beijing is located in the temperate zone and enjoys a continental monsoon climate. Late August is still quite hot in the daytime but cool at night, with an average of 24.5 degrees Celsius or 76 degrees Fahrenheit; the high temperature in the day can reach 33 C or 91 F and the low could be 18 C or 65 F. There is not much rain during this time. So travel light to Beijing. A jacket will be sufficient for the cool rainy day, and an umbrella is necessary for both scorching sun and sudden rain if you are going out.

H.4 Electric Current

The electric current is 220 V (50 Hz). The hotels can provide 220v and 110v (shavers only) power outlets. Please note that adapters may be necessary.

H.5 Bank Services

RMB (Renminbi) is the only currency used in China. Money exchanges by cash or traveler’s cheques can be made at the branches of Bank of China at Beijing International Airport, hotels and tourist stores. There is also a bank in the basement of the BICC. Banks usually open at 8:00-9:00 in the morning and close at 17:00 in the afternoon everyday of the week. Conversion services are available for the following foreign currencies: US Dollar, Euro, British Pound Sterling, Japanese Yen, Hong Kong Dollar, Australian Dollar, Canadian Dollar, Swiss Franc, Singapore Dollar, Danish Krone, Norwegian Krone, Swedish Krone, Malaysian Ringgit and Macau Pataca. But note that not all these currencies are accepted at some small money exchange outlets. In the basement of the BICC, the Congress venue, China Merchant Bank provides money exchange service for US Dollar, Euro, British Pound Sterling, Japanese Yen, Hong Kong Dollar, Australian Dollar, Canadian Dollar, Swiss Franc and Singapore Dollar. The opening hours are 8:30-18:30 every day of the week. In all star-rated hotels, there are also money exchange services.

H.6 Credit Cards

Credit cards such as American Express, Diners Club, JCB, Master and Visa are acceptable for payment at hotels or tourist stores. But it is difficult to draw cash with credit cards. It would be best to change money with internationally recognized traveler’s checks, or foreign cash.

H.7 Shopping Hours

Opening hours are 8:30-21:00 for major department stores, every day of the week.

H.8 First Aid, Health and Accident Insurances

The congress fee does not include insurance for the participants against accidents, sickness, or loss of personal property. All participants are strongly advised to make necessary arrangements for short–term health and accident insurance in advance. In any case, the organizers refuse all liability to cover health or accident expenses of participants unless expenses are due to an act of negligence by ICM2002.

During the congress, First Aid will be available in the BICC. In case of emergency, please contact the registration counter, the congress office, or one of the lecture room attendees. In case of illness, you may go to the hospitals or clinics listed below.

• Beijing Union Hospital, Tel. 65296114

• International SOS Assistance, Tel. 64629112, 64629100, 14 Liangma He South Road

• Sino-Japanese Friendship Hospital, Tel. 64221122

• International Medical Center, Tel. 64651561/2/3, S106 Lufthansa Business Center

• AEA International, Tel. 64651384, S103 Lufthansa Business Center

• Hong Kong International Clinic, Tel. 65122288 ext. 2346, 3/F, Swiss Hotel

I. Registration

I.1 Membership

It has been a long tradition to call any person who has registered for ICM2002 an Ordinary Member of the Congress. Registration is required in order to be admitted to the scientific program of the Congress. Ordinary members will receive a registration package including a congress badge, the program, an abstract book, and other material at the registration counter as well as a copy of the Proceedings when published. Furthermore the registration fee includes the Opening Ceremony with the Opening Reception, coffee during coffee breaks, and the ICM2002 Party.

Please do wear your congress badge at all congress activities or whenever you want to be recognized as a congress member; in any case, be prepared to show the badge when asked to do so.

Students not having completed their PhD may register at the reduced student rate by supplying an official student certificate of their university. Student registration also includes the Proceedings volume, thanks to the supports from various ministries of the Chinese government.

I.2 Accompanying Persons

Only registered accompanying persons are entitled to attending the Opening Ceremony, Opening Reception after the Opening Session, the ICM2002 Party, and coffee breaks.

I.3 Registration

There are two possible ways for ICM2002 registration. You may register electronically at the following registration site on line: http://www.icm2002.org.cn/I/Register.htm. Alternatively, you may complete the attached registration form and return it to the CICCST with the address provided in F.2.

I.4 Registration Fees

Registration  until  after
April 30, 2002 April 30, 2002
Full registration fee  US$ 240 ,–  US$ 280,–
Students*  US$ 120,–  US$ 140,–
Accompanying Persons†  US$ 80,–  US$ 100,–

The full registration or student registration fee includes all the conference materials and the Proceedings of the Congress, the Opening Reception after the Opening Session, and the ICM2002 Party.

* Registration as a student requires an official certificate of the university to be attached. † Registered accompanying persons are entitled to attend the Opening Session, Opening Re-

ception after the Opening Session, the ICM2002 Party, and coffee breaks.

I.5 Methods of Payment

All fees must be paid in U.S. dollars, by bank transfer, bank draft or credit card. Currency exchange charges or bank service fees are the money sender’s responsibility. If these are not paid or have occurred during money transfer, they will be charged upon registration at the registration desk.

1. By bank transfer: Account No.: 00038018241014, Account Name: CICCST, Bank Name: Bank of China Head Office, 1 Fu Xing Men Nai Da Jie, Beijing 100818, China. A duplicate or photocopy of the remittance order from the bank should be sent with the Registration Form to the Congress Secretariat. The name of the participant as well as details of payment should be clearly marked.

2. By Bank Draft, made payable to CICCST/ICM2002, mail to Mr. Liu Feng, ICM2002, CICCST, 86 Xueyuan Nan Road, Beijing 100081, China.

3. By Credit Card: Fill out the appropriate place in the registration form and sign your name, or, if you register on-line, download a copy of the Payment Authorization Form from

     http://www.icm2002.org.cn/I/Registration.htm, and fill it out and fax it to Mr. Liu

Feng, +86-10-6218 0142, or mail to him at the address given in section F.2.

Note that only American Express, Diners Club, JCB, Master and Visa are acceptable. And, as the payments are calculated in Chinese RMB, the amount charged to your account may vary slightly from the amount billed, due to fluctuations in exchange rates. On-site payment should be made in US dollars in cash or traveler’s checks or credit card.

I.6 Cancellations

All cancellations of registration, hotel reservation, social events and tours booking must be in written form. No refund of registration will be made after July 15, 2002 but substitute delegates are acceptable. For pre- or post-Congress tours, no refund will be made if the cancellation is received five days before the start of the tour. Refunds will be made from September 1, 2002 where applicable.

Cancellation Received  Before July 15, 2002 From July 16, 2002 After August 10, 2002
Delegate  US$150 refunded  No Refund
Student  US$ 80 refunded  No Refund
Accompanying Person  US$ 60 refunded  No Refund
Hotel reservation  80% refunded  60% refunded  No Refund

Local tours, tickets

for Social events  80% refunded  60% refunded  No Refund
Pre/Post Congress tours  90% refunded  60% refunded  see above

I.7 Registration Counter

The registration counter will be set up in the lobby of Beijing International Convention Center and open in the following hours:

      Monday       August 19, 2002          9:00-21:00
      Tuesday      August 20, 2002          8:00-21:00
      Wednesday    August 21, 2002          8:00-18:00
      Thursday     August 22, 2002          8:30-18:00
      Friday       August 23, 2002          8:30-18:00
      Saturday     August 24, 2002          8:30-18:00
      Sunday       August 25, 2002          8:30-18:00
      Monday       August 26, 2002          8:30-18:00
      Tuesday      August 27, 2002          8:30-18:00
      Wednesday    August 28, 2002          8:30-12:00

If fees have been forwarded late and therefore not yet been credited to the account of CICCST on the day of arrival, a copy of the remittance order should be presented.

J. Accommodation

The CICCST has blocked as many as possible hotel rooms for intending participants around the Congress venue, BICC. Some of these hotels are only 15 to 20 minutes walk from the BICC. The others are only 20-30 minutes by shuttle bus, which will be provided by the CICCST. All these hotels are air-conditioned. If you wish to make hotel reservation for the Congress via CICCST, please use the attached registration form or the internet on-line registration. Reservation of rooms will be made in accordance with your first choice, or the second choice if the rooms of your first choice have been booked out. Once the requested reservation has been made, you will receive confirmation via email.

J.1 Hotels

A block of rooms will be reserved for Congress participants and accompanying persons in the hotel categories listed below at special rates for standard two-bed rooms per night. These hotels are all air-conditioned. Please note that hotel confirmations are subject to a deposit of US$100 (which will be deducted from your final hotel bill). Congress participants are requested to use the Registration Form or the on-line registration for the hotel reservation. They are advised to check at the Congress registration website http://www.icm2002.org.cn whether the rooms of their choice are still available before they make a reservation on-line. Reservation should be made before June 15, 2002. Reservation later than this date may not be guaranteed of a room of your choice.

Category A: US$55-90 Category B: US$45-55 Category C: US$35-45 Category D: US$25-35

A few of the hotels of Category A are located near the Congress venue, such as Beijing Continental Grand Hotel and CATIC Hotel.

J.2 Apartments

A block of apartment rooms will be reserved for Congress participants and accompanying persons. These apartments are located within a 10 minutes walk from the Congress venue. These apartments have two, three or four bedrooms, each with two or three beds. The occupants of the bedrooms will have to share common toilets in the apartment. Every apartment is furnished with basic furniture, a TV set, and a telephone for local calls. These apartments are very suitable for groups of several people. The cost is around US$13-16 per bed per night. In your reservation, indicate how many beds in how many rooms you want to reserve. To reserve the apartments, follow the procedure for the reservation for hotel rooms, but the balance due will be paid to the Congress Secretariat directly. The minimal reservation will be two beds in one room. The deadline for the reservation of apartments is May 15, 2002.

J.3 Student Dormitories

Student dormitories in universities might be available to participants during ICM2002. But no actual information can be obtained until March 1, 2002. Participants interested in reservation is requested to inquire or visit the Congress website for latest information after March 1, 2002.

J.4 Liability

In all cases the organizers of ICM2002 shall only act as an agent and shall not be liable for any losses, accidents, personal injury, or damage to property of any kind and origin whatsoever. The liability of the persons or companies instructed to provide the services shall not be affected. Amendments shall not be valid unless mutually agreed in writing and signed by the two contracting parties. The place of jurisdiction for fully qualified merchants, persons who do not have a general place of jurisdiction in China, and persons who have relocated their place of residence to a foreign country after the conclusion of the agreement shall be Beijing.

K. Social and Tourist Program

Overview

Time  Price in US$

Tuesday August 20, 2002 Opening Ceremony in the Great Hall of the People

followed by a Opening Reception  15:00–20:30 free of charge*

Wednesday August 21, 2002

The Great Wall at Juyong Pass and the Ming Tombs  8:30–17:00  38
The Forbidden City and the Temple of Heaven  8:30–17:00  32
Museum of Chinese History, the Zoo and the Silk Street 8:30-16:30  24

Thursday August 22, 2002

The Great Wall at Juyong Pass and the Ming Tombs  8:30–17:00  38
Summer Palace and the Lama Temple  8:30–16:30  26
Tiananmen Square, Beihai Park, and the Lane Tour  8:30–16:30  26

Friday August 23, 2002

The Great Wall at Juyong Pass and the Ming Tombs  8:30–17:00  38
Summer Palace and the Lama Temple  8:30–16:30  26
Museum of Chinese History, the Zoo and the Silk Street 8:30-16:30  24

Saturday August 24, 2002

The Great Wall at Juyong Pass and the Ming Tombs  8:30–17:00  38
The Forbidden City and the Temple of Heaven  8:30–17:00  32
Tiananmen Square, Beihai Park, and the Lane Tour  8:30–16:30  26

Sunday August 25, 2002

The Forbidden City and the Temple of Heaven  8:30–17:00  32
Summer Palace and the Lama Temple  8:30–16:30  26
Museum of Chinese History, the Zoo and the Silk Street 8:30-16:30  24
Beijing Opera 19:30–20:45 10-24

Monday August 26, 2002

The Great Wall at Juyong Pass and the Ming Tombs  8:30–17:00  38
The Forbidden City and the Temple of Heaven  8:30–17:00  32
Tiananmen Square, Beihai Park, and the Lane Tour  8:30–16:30  26
ICM Party  19:00-21:00 free of charge†

Tuesday August 27, 2002

The Great Wall at Juyong Pass and the Ming Tombs  8:30–17:00  38
Summer Palace and the Lama Temple  8:30–16:30  26
Museum of Chinese History, the Zoo and the Silk Street 8:30-16:30  24

All bus tours start in front of the Beijing International Convention Center. Please note that the return times of the bus tours are approximate and depend on traffic conditions. Luncheons are included beverages. The program is subject to change. All tours are provided with English speaking guides.

* Only for participants and registered accompanying persons. † For un-registered accompanying persons, each ticket costs $20.

K.1 Social Events

Opening Reception Date: August 20, 2002 Time: 19:00 after the Opening Session Place: The Great Hall of the People Transportation: Provided by the CICCST, at 14:00 from the Congress-designated hotels ICM2002 Party Date: August 26, 2002 Time: 19:00 Place: The lawn in front of the BICC

Beijing Opera Date: August 25, 2002 Time: 19:30 Place: Chang’an Theatre Transportation: Provided by the CICCST. Buses start in front of the BICC. Ticket Categories: First Category US$ 24, Second Category US$ 16, Third Category US$ 13, Fourth Category US$ 10. If tickets in the desired category are sold out, tickets in another category will be reserved. Following is a description of the three selections of the Opera.

1. The Crossroads: Jiao Zan, a senior officer of the Song Dynasty, is being exiled under guard to Shamen Island because he has killed a treacherous court official. Marshel Yang orders Ren Tanghui to protect Jia in secret. Ren and Jiao put up for the night at the Crossroad Inn. The inn-keeper Liu Lihua believes that Ren intended to murder Jiao, so he steals into Ren’s room and fight with him in the dark. Only when the inn-keeper’s wife comes in with a candle do the three realize in the end that all has been the result of misunderstanding.

2. Stealing Magic Herbs: This is an episode of The Romance of the White Snake. On the day of Dragon Boat Festival, Xu Xian advises his wife, Bai Suzhen, for a drink of medicated wine. At her dead drunkenness, Bai shows herself in her true colors - a white snake. At the sight of this, Xu is scared to death. To save her husband, Bai goes to the forbidden mountains to steal magic herbs. There she fights with the guards - crane boys and deer boys, and gets the right herbs she wants.

3. Farewell My Concubine (The Death of Yu Ji): Liu Bang and Xiang Yu have agreed to a truce and drawn a demarcation line at Honggou. Liu’s general makes a feign surrender to Xiang and then successfully lures Xiang and his troops into an ambush. Xiang and his troops are surrounded and cannot break through. When his soldiers hear their folk songs sung from the enemy, they take it for granted that their fellow soldiers have given up fighting, and their morale goes down. Xiang realizes that the game is as good as lost and indulges in his drinks in despair. He bids farewell to his lover, Yu Ji, who dances her last before killing herself with a sword.

K.2 Tourist Program

All tours, either local or pre- and post-Congress, are operated by China Travel Service Head Office. To make reservations, fill out the section of Tours Reservation in the Registration Form.

Local Tours

As the ancient capital of five dynasties - the Liao, Jin, Yuan, Ming and Qing - Beijing boasts numerous relics and beautiful historical monuments such as the Great Wall, the magnificent imperial palace, different religious temples and imperial gardens. The local tours will introduce you to the old and new Beijing. All the local tours are one-day tours and start and end in front of the BICC. LT-1: The Great Wall at Juyong Pass and the Ming Tombs During this one-day tour, you will visit a section of the Great Wall at Juyong Pass and one of the Ming Tombs. The Great Wall stretches 6000 kilometers from East to West China, traversing five provinces and two autonomous regions. Built first in 221 B.C. and re-built several times in the following 1500 years, it has become a symbol of the ancient Chinese civilization. After the visit to the Wall and lunch, you will visit an emperor’s tomb of the Ming Dynasty (1380-1644). This tomb is a mausoleum of the 13th Ming emperor and is the only one that has been excavated. And, the best part of the tomb area is the Spirit Way – giant stone carvings of human beings, animals and mythical monsters standing guard. US$ 38 per person, lunch included. Tour operated on

         Aug. 21  8:30-17:00           Aug. 22  8:30-17:00
         Aug. 23  8:30-17:00           Aug. 24  8:30-17:00
         Aug. 26  8:30-17:00           Aug. 27  8:30-17:00

LT-2: The Forbidden City and the Temple of Heaven In this one-day tour you will see the imperial palace (officially called the Palace Museum) and the Temple of Heaven. The Forbidden City was constructed in 1406. It was the imperial palace of the Ming and Qing Dynasties from 1420 to 1911. An architectural masterpiece, it was China’s political center for 500 years. It is now converted into a museum with a large imperial collection of porcelain, goldware, bronzeware, jadeware, traditional paintings, and jewelry. After lunch, you will visit the Temple of Heaven. The Temle of Heave was built 600 years ago. It served as a means of communication between the Emperor and the Heaven. The buildings are a masterpiece of ancient architecture. The focal points are the Hall of Prayer for Good Harvests (a completely wooden structure), the Imperial Vault of Heaven (with a circular Echo Wall with acoustic effects), and the Circular Mound Altar (designed with geometric ingenuity). US$ 32 per person, lunch included. Tour operated on:

         Aug. 21  8:30-17:00          Aug. 24  8:30-17:00
         Aug. 25  8:30-17:00          Aug. 26  8:30-17:00

LT-3: Summer Palace and the Lama Temple In the morning, you will visit the Summer Palace, the largest and best preserved imperial garden. It used to be a retreat for emperors to escape the scorching summer heat in Beijing. The garden is so large that it is scarcely enough to see in one day all the halls, towers, pavilions, corridors, walkways and bridges that grace the hill slopes, lake shore and islets. In the afternoon, you will visit the Lama Temple. The temple is the most impressive and an active Buddhist temple in the city. One of the prominent features of the temple is the 23-meter-tall Maitreya Buddha, which was carved out of a whole trunk of white sandalwood tree from Tibet. US$ 26 per person, lunch included. Tour operated on:

         Aug. 22  8:30-16:30         Aug. 23  8:30-16:30
         Aug. 25  8:30-16:30         Aug. 27  8:30-16:30

LT-4: Tiananmen Square, Beihai Park, and the Lane Tour In the morning you will visit the Tiananmen Square and Beihai Park. Tiananmen Square is located in the center of the city, to the south of the Forbidden City. It is the largest city square in the world. At the Square there are Mao’s Mausoleum, the Museum of Chinese History and the Great Hall of the People (parliament building). Beihai Park is to the northwest of the Forbidden City. It was built on the basis of a legend and used to be a royal garden. After lunch, you will visit the backyards of Beijing, on a pedicab, across the small zig-zag lanes. US$ 26 per person, lunch included. Tour operated on:

         Aug. 22  8:30-16:30          Aug. 24  8:30-16:30
         Aug. 26  8:30-16:30

LT-5: Museum of Chinese History, the Zoo and the Silk Street In the morning you will visit the Museum of Chinese History where you will see various archaeological findings that illustrate the development of the Chinese nation. After lunch, you will visit the Beijing Zoo and see the Giant Panda, an endangered species of animal native to China. Then, you will have a walk around the Silk Alley, a shopping street (market) in the embassy area. US$ 24 per person, lunch included. Tour operated on:

        Aug. 21  8:30-16:30         Aug. 23  8:30-16:30
        Aug. 25  8:30-16:30         Aug. 27  8:30-16:30

Pre- and Post-Congress Tours

The pre- and post-Congress tours will show you the most beautiful parts of China. Reservation of a pre-Congress tour is subject to the payment of full cost by May 31, 2002, and reservation of a post-Congress tour is subject to a deposit of US$300 by May 31, 2002. Any one these tours will take place when there are a minimum of ten participants for the tour. If there are insufficient bookings for a tour, full refund will be made for this tour.

The price includes domestic inter-city air or rail transportation, accommodation in three-star hotels or hotels of the same level, regular meals, admission tickets for the sites specified, and local transportation. It does not include airport tax, air-trip accident insurance, overweight of luggage, personal expenses, and tips.

For the detailed itinerary, please visit http://www.icm2002.org.cn. PRE-1: Beijing-Yichang-Chongqing-Beijing August 14-19, 2002 Twin Occupancy: US$666; Single Supplement (only in Beijing): US$25 This trip features an upstream boat cruise through the Three Gorges along the world’s third longest river, the Yangtze River. You will first visit in Yichang the Chinese Sturgeon, an endangered species of fish, and the Gezhou Dam on the Yangtze. Then, you will get on the boat for the three-day cruise. The boat arrives at Chongqing, the capital city of China during the World War II, and you will fly back to Beijing from there. PRE-2: Beijing-Urumqi-Turpan-Dunhuang-Xi’an-Beijing August 11-19, 2002 Twin Occupancy: US$1099; Single Supplement: US$180 This Silk Road Tour will introduce you to the Northwestern China. You will first fly from Beijing to Urumqi, the capital city of Xinjiang and a scenic spot along the ancient Silk Road. Then you will visit several famous places along the road, like the Old City Between Two Rivers (Jiaohe Gucheng) in Turpan, Mogao Grottoes in Dunhuang, the underground army of Terra Cotta Warriors and Horses in Xian. Coach ride and overnight train ride will be used during the tour. PRE-3: Shanghai-Guilin-Xi’an-Beijing August 14-19, 2002 Twin Occupancy: US$630; Single Supplement: US$75 This tour is a combination of landscape beauty and archeological grandeur. When you arrive in Shanghai, China’s industrial and financial center, you will take a tour to the Bund, and the Yu Garden. Then, you will fly to Guilin to have a half day boat cruise on the Li River and visit some sites of karst landscape and caverns, the most spectacular in China. From Guilin you will fly to Xian, which boats China’s best archeological findings, to visit the underground army of Terra Cotta Warriors and Horses and Banpo Museum of remains of a settlement in the Neolithic age. The tour will end in Beijing. PRE-4: Kunming-Guilin-Xi’an-Beijing August 12-19, 2002 Twin Occupancy: US$640; Single Supplement: US$95 This tour is similar to PRE-3. Instead of Shanghai, you should arrive in Kunming, capital of the southwestern province Yunnan to meet the group. There you will have a day trip to the Stone Forest and Western Hills. The other sites in Guilin and Xian are the same as in PRE-3. PT-1: Beijing-Xi’an-Beijing August 29-30, 2002 Twin Occupancy: US$335; Single Supplement: US$15 This is a short two-day tour for those who do not have sufficient time to see more of China. During this trip you will visit the underground army of Terra Cotta Warriors and Horses, Banpo Museum of remains of a settlement in the Neolithic age, and the Provincial Museum of History.

PT-2: Beijing-Xi’an-Guilin August 29-September 2, 2002 Twin Occupancy: US$455; Single Supplement: US$66 This tour is similar to the pre-Congress tour PRE-3, but without Shanghai. You will visit the underground army of Terra Cotta Warriors and Horses and Banpo Museum in Xian, and have a boat cruise along the Li River in Guilin. The tour will end in Guilin, and you can fly out to Hong Kong or other Chinese cities by your own arrangement. PT-3: Beijing-Xi’an-Chengdu-Kunming August 29-September 5, 2002 Twin Occupancy: US$560; Single Supplement: US$95 During this trip, you will first visit the underground army of Terra Cotta Warriors and Horses, Banpo Museum of remains of a settlement in the Neolithic age, and the Provincial Museum of History in Xian. Then you will fly to Chengdu, an important city in Southwestern China, to visit the Panda Research Center and Dujiang Dam, an ancient water conservancy project. In Kunming you will visit the Western Hills and take a day trip tour to the Stone Forest. The tour will end in Kunming, and you can fly out to Hong Kong or other Chinese cities by your own arrangement. PT-4: Beijing-Xi’an-Chengdu-Lhasa-Beijing August 29-September 5, 2002 Twin Occupancy: US$1030; Single Supplement: US$130 In this tour, in addition to the sites of Xian and Chengdu as listed in PT-3, you will visit Lhasa, the political and economic center of Tibet for the last 360 years. There you will be shown around to tour the Potala Palace, Drepang Temple and Norbu Linka. High-altitude reaction may be experienced during the tour in this city. PT-5: Beijing-Chongqing-Yichang-Shanghai August 29-September 5, 2002 Twin Occupancy: US$740; Single Supplement: US$65 This is a two-day downstream boat cruise through the Three Gorges along the world’s third longest river, the Yangtze. After the cruise, you will visit the Chinese Sturgeon, and the Gezhou Dam on the Yangtze in Yichang. Then you will fly to Shanghai, to tour the Bund, the Jade Buddha Temple and the Yu Garden. The tour will end in Shanghai. PT-6: Beijing-Nanjing-Suzhou-Hangzhou-Shanghai August 29-September 4, 2002 Twin Occupancy: US$330; Single Supplement: US$85 This trip will show you the eastern part of China, which is claimed to be the paradise on earth. You will first fly to Nanjing, capital of the Nationalist government before 1949. You will visit the Mausoleum of Dr. Sun Yat-sen, founder of the Republic, and the Yangtze River Bridge. Then you will have a coach ride to Suzhou to tour the typical Chinese-style gardens. Then the coach will drive you to Hangzhou to see the Lingyin Temple and West Lake. The coach tour will end in Shanghai, where you will visit the Bund, the Jade Buddha Temple and the Yu Garden. The tour will end in Shanghai.

L. Satellite Conferences

Listed in the sequence of the date when the proposal was received.

International Congress of Mathematical Software. Beijing. Aug. 17-19. Contact:

 xgao@mmrc.iss.ac.cn; zliu@mmrc.iss.ac.cn. Website: http://www.mathsoftware.org/

Operator Algebras and Applications.  Chengde.  Aug.14-18.  Contact:

 chengde@OAinChina.org; hxlin@cartan.uoregon.edu. Website: http://www.OAinChina.org

Differential Geometry and Global Analysis. Tianjin. Aug. 17-18. Contact:

 weiping@sun.nankai.edu.cn

Mathematical Biology. Guilin. Aug. 15-18. Contact: gxnu@public.glptt.gx.cn; waiban@mailbox.gxnu.edu.cn. Website: http://math.la.asu.edu/˜ kuang/guilin.html

Symplectic Topology and Geometry. Chengdu. Aug. 14-18. Contact: gszhao@scu.edu.cn. Website: http://www.scu.edu.cn

Ordinary Differential Equations. Lanzhou (Dunhuang). Aug. 30-Sept. 4. Contact: antianqing@21cn.com. Website: http://www.nwnu.edu.cn

Discrete, Combinatorial and Computational Geometry. Beijing. Aug. 13-19. Contact:

cmzong@math.pku.edu.cn

Combinatorics. Shijiazhuang. Aug.30-Sept. 3. Contact: qdkang@public.sj.he.cn; rending@public.sj.he.cn. Website: http://math.hebtu.edu.cn

Nonlinear Functional Analysis. Taiyuan. Aug. 11-19. Contact: sjli@mail.sxu.edu.cn. Website: http://www.math.sxu.edu.cn

International Colloquium for the History of Mathematics. Xi’an. Aug. 15-18. Contact:

hismath@nwu.edu.cn; qaj@sein.sxgb.com.cn

Mathematics Education. Lhasa. Aug. 12-17. Contact: xzsxh@utibet.edu.cn

New Directions in Dynamical Systems.  Kyoto.  Aug.  5-15.  Contact:

kokubu@kusm.kyoto-u.ac.jp. Website: http://ndds.math.h.kyoto-u.ac.jp/index.html

Matrix Theory and Its Application.  Shanghai.  Aug.  14-18.  Contact:

ejiang@fudan.edu.cn; chyu-k@online.sh.cn

Scientific Computing. Xi’an. Aug. 15-18. Contact: ktli@xjtu.edu.cn. Website: http://www.xjtu.edu.cn/xjnet/temp/icm2002/shouye.htm

Clifford Analysis. Macau. Aug. 16-19. Contact: fsttq@umac.mo. Website: http://www.umac.mo/˜ Clifford2002/

Differential and Functional Differential Equations. Moscow. Aug. 11-17. Contact:

k803@mai.ru

Control Theory and Optimization.  Xi’an.  Aug.  30-Sept.  1.  Contact:

quanpan@nwpu.edu.cn; tanggp@nwpu.ac.cn. Website: http://www.nwpu.edu.cn/ICM-2002-Xi’an

Algebra. Suzhou. Aug. 29 - Sept. 2. Contact: yysx@suda.edu.cn; zmtang@suda.edu.cn. Website: http://math.suda.edu.cn/icm/ALGEBRAIC/topics.html

Nonlinear Evolution Equations and Dynamical Systems. Huangshan. Aug. 29-Sept. 1.

Contact: chengy@ustc.edu.cn

Algebraic Geometry. Shanghai. Aug. 13-17. Contact: ag2002@wims.math.ecnu.edu.cn. Website: http://wims.math.ecnu.edu.cn/ag2002

Computational Mathematics and Applications. Dalian. Aug. 30-Sept. 3. Contact:

renhong@dlut.edu.cn; mateokl@polyu.edu.hk; iscm02@dlut.edu.cn. Website: http://

www.acad.polyu.edu.hk/˜ math; http://www.dlut.edu.cn/dutn/index pages/iscma02/index.htm

Geometric Function Theory in Several Complex Variables. Hefei. Aug. 30 - Sept. 2. Contact: lts@ustc.edu.cn. Website: http://www.ustc.edu.cn/ICM2002GFT Nonlinear Partial Differential Equations-Theory and Approximation. Hong Kong. Aug. 29-Sept. 2. Contact: malam@cityu.edu.hk

Number Theory and Arithmetic Geometry. Weihai. Aug. 13-17. Contact: icm2002nt@sdu.edu.cn. Website: http://www.sdu.edu.cn

Nonlinear Partial Differential Equations in Mechanics and Physics. Harbin-Mudanjiang(Jingbo Lake). Aug. 29- Sept.3. Contact: sxx zr@hrbnu.edu.cn

Harmonic Analysis and its Applications. Hangzhou. Aug. 14-18. Contact:

jcchen@mail.hz.zj.cn

Backward Stochastic Differential Equations. Weihai and Beijing. Aug. 29-31. Contact:

dice@math.sdu.edu.cn

International Conference on Mathematical Logic Chongqing. Aug. 29-Sept. 2. Contact:ytang@swnu.edu.cn. Website: http://www.swnu.edu.cn

Symposium on Stochastics and Applications. Singapore, Aug. 15-17. Contact: ssa@math.nus.edu.sg. Website: http://www.math.nus.edu.sg/ssa

Geometric Topology.  Xi’an.  Aug.  12-16.  Contact: xl@math.ucr.edu;

wangsc@math.pku.edu.cn. Website: http://www.math.uiowa.edu/˜ wu/gtc/gtc.html

Game Theory and Applications.  Qingdao.  Aug.  14-17.  Contact:

gaosai@public.qd.sd.cn; spbuoasis7@peterlink.ru. Website: http://www.mathqdu.org;

http://isdgrus.apmath.spbu.ru

Combinatorics, Graph Theory and Applications. Hong Kong. Aug.15-17. Contact: mabfchen@ust.hk. Website: http://www.math.ust.hk/conference

Infinite Dimensional Function Theory.  Pohang.  Aug.  12-16.  Contact:

mathchoi@euclid.postech.ac.kr

Several Complex Variables and Complex Geometry. Daejeon. Aug. 14-18. Contact:

kimkt@postech.edu; kimkt@math.wustl.edu. Website: http://math.postech.ac.kr

International Conference in Algebras and Related Topics. Hong Kong. Aug. 14-17.

Contact: KPShum@math.cuhk.edu.hk

Complex Analysis. Shanghai. Aug. 14-17. Contact: anfang@mail.sjtu.edu.cn;

wangyf@math03.math.ac.cn

Electronic Information and Communication in Mathematics. Beijing. Aug.

29-31. Contact: icms@math.tsinghua.edu.cn; nsu@math.tsinghua.edu.cn. Website:

http://icms.math.tsinghua.edu.cn

The 7th International Conference on Difference Equations and Applications. Changsha, Aug, 11-17. Contact: applmath@mail.hunu.edu.cn. Website: http://www.math.mun.ca/˜ xzou/7thICDEA/main.html

Algebraic Topology. Suzhou. Aug. 30-Sept. 3. Contact: yuyl@suda.edu.cn. Website: http://math.suda.edu.cn/icm/topology/atc.html

Bifurcation and Chaos. Kunming. Aug. 13-17. Contact: jibinli@hotmail.com;

huifang@public.km.yn.cn

International Conference on the Reform of Mathematics Curriculum and its Education in the 21st Century. Chongqing. Aug. 17-19. Contact: math@swnu.edu.cn. Website: http://www.swnu.edu.cn/

Nonlinear Analysis. Taipei. Aug. 13-17. Contact: icm2002@math.sinica.edu.tw. Website: http://math.sinica.edu.tw/icm2002

Abstract and Applied Analysis. Hanoi. Aug. 13-17. Contact: nmchuong@

thevinh.ncst.ac.vn; nhdien@thevinh.ncst.ac.vn; huongnga@thevinh.ncst.ac.vn

Stochastic Analysis. Beijing. Aug. 29 - Sept. 3. Contact: Wang@uni-bonn.de;

dayue@math.pku.edu.cn

Fractal Geometry and Applications. Nanjing. Aug. 30-Sept. 2. Contact:

suqiu@netra.nju.edu.cn; wenzy@tsinghua.edu.cn

String Theory Conference. Hangzhou. Aug. 12-15. Contact: Xuhongwei@math.hz.zj.cn. Website: http://www.zju.edu.cn/string2002

M. ICM2002 Sponsors

The ICM2002 Organizing Committee has maintained the tradition of setting a relatively low registration fee to attract as more mathematicians as possible. Nevertheless, there will still be many mathematicians for whom participation is economically unaffordable. To help at least some of them to attend, ICM2002 Organizing Committee has, in cooperation with the International Mathematical Union, set up a program to offer financial support for active young mathematicians as well as mature mathematicians from developing countries. Details can be found at the ICM2002 Web Server http://www.icm2002.org.cn. With all this, the registration fee only cannot cover all costs of ICM2002. The various ministries of the Chinese government has rendered strong support without which the success of ICM2002 cannot be guaranteed. The Organizing Committee of ICM2002 are very grateful to the following:

• Ministry of Finance of PRC

• Ministry of Science and Technology of PRC

• Ministry of Education of PRC

• Chinese Academy of Sciences

• National Natural Science Foundation of China

• Beijing Municipal Government

• China Association for Science and Technology

Finally we would like to thank Shanghai Municipal Government, who will sponsor the IMU General Assembly.

The IMU General Assembly will take place in Shanghai, China, on August 17-18, 2002.

N. 2000 Mathematics Subject Classification

00 General 01 History and biography 03 Mathematical logic and foundations 05 Combinatorics 06 Order, lattices, ordered algebraic structures 08 General algebraic systems 11 Number theory 12 Field theory and polynomials 13 Commutative rings and algebras 14 Algebraic geometry 15 Linear and multilinear algebra; matrix theory 16 Associative rings and algebras 17 Nonassociative rings and algebras 18 Category theory, homological algebra 19 K-theory 20 Group theory and generalizations 22 Topological groups, Lie groups 26 Real functions 28 Measure and integration 30 Functions of a complex variable 31 Potential theory 32 Several complex variables and analytic spaces 33 Special functions 34 Ordinary differential equations 35 Partial differential equations 37 Dynamical systems and ergodic theory 39 Difference and functional equations 40 Sequences, series, summability 41 Approximations and expansions 42 Fourier analysis 43 Abstract harmonic analysis 44 Integral transforms, operational calculus 45 Integral equations 46 Functional analysis 47 Operator theory 49 Calculus of variations and optimal control; optimization 51 Geometry 52 Convex and discrete geometry 53 Differential geometry 54 General topology 55 Algebraic topology 57 Manifolds and cell complexes 58 Global analysis, analysis on manifolds 60 Probability theory and stochastic processes 62 Statistics 65 Numerical analysis 68 Computer science 70 Mechanics of particles and systems 74 Mechanics of deformable solids 76 Fluid mechanics 78 Optics, electromagnetic theory 80 Classical thermodynamics, heat transfer 81 Quantum Theory 82 Statistical mechanics, structure of matter 83 Relativity and gravitational theory 85 Astronomy and astrophysics 86 Geophysics 90 Operations research, mathematical programming 91 Game theory, economics, social and behavioral sciences 92 Biology and other natural sciences 93 Systems theory; control 94 Information and communication, circuits 97 Mathematics education

O. Organizing Committee

The Organizing Committee was set up in 1998 and has been growing since. At the moment of this writing the following people are involved:

CHAIRMAN Ma, Zhi-Ming, Inst of Appl Math, AMSS, CAS, Beijing.

MEMBERS Chang, Kung Ching, Peking University, Beijing; Chen, Shuping , Zhejiang University, Hangzhou; Ding, Weiyue, Inst of Math, AMSS, CAS, Beijing; Feng, Keqin, Tsinghua University, Beijing; Feng, Qi, Inst of Math, AMSS, CAS, Beijing; Hou, Zixin, Nankai University, Tianjin; Jiang, Boju, Peking University, Beijing; Li, Daqian, Fudan University, Shanghai; Li, Wenlin, Inst of Math, AMSS, CAS, Beijing; Lin, Qun, Inst of Comp Math, AMSS, CAS, Beijing; Lin, Fanghua, University of New York, New York; Liu, Taiping, Stanford University, San Francisco; Liu, Yingming, Sichuan University, Chengdu; Lu, Shanzhen, Beijing Normal University, Beijing; Ma, Zhi-Ming, Inst of Appl Math, AMSS, CAS, Beijing; Peng, Lizhong, Peking University, Beijing; Wang, Jianpan, Eastern China Normal University, Shanghai; Wong, Roderick, City University of Hong Kong, Hong Kong; Yang, Lo, Inst of Math, AMSS, CAS, Beijing; Yuan, Ya-xiang, Inst of Comp Math, AMSS, CAS, Beijing; Zhang, Jiping, Peking University, Beijing; Zhang, Xiangsun, Inst of Appl Math, AMSS, CAS, Beijing; Zhou, Qing, Eastern China Normal University, Shanghai.

28 29